OpenStudy (gg):

Solve ordinary differential equation xydy=(y^2+x)dx Solve ordinary differential equation xydy=(y^2+x)dx @Mathematics

OpenStudy (amistre64):

D[xy dy = (y^2+x)dx] y = 2y y' well, its not an exact is it

OpenStudy (gg):

u don't have to write it all. I got udu=(1/x^2)dx (u=y/x). what now?

OpenStudy (amistre64):

i spose the ^2 makes this nonlinear?

OpenStudy (amistre64):

they aint homogenous either are they

OpenStudy (amistre64):

im still starting these so I aint no expert yet ;)

OpenStudy (amistre64):

you done bernoullis yet?

OpenStudy (gg):

yes, but this is not bernoullis

OpenStudy (amistre64):

yeah, i didnt think so either; but it would have been nice if it was lol xy dy = (y^2+x) dx xy y' = y^2+x y' = (y^2+x)/xy y' = y/x +1/y hmmmm

OpenStudy (amistre64):

y' - (1/x)y = y^(-1) you sure we cant do a bernie on it?

OpenStudy (gg):

well, I suppose that there's not just one way to solve this :)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%27+%3D+y%2Fx+%2B+1%2Fy i got some idea, but its prolly not correct

OpenStudy (gg):

hahha, I used wolphram too :D

OpenStudy (gg):

*wolfram

OpenStudy (amistre64):

yeah, a the moment im at a loss

OpenStudy (amistre64):

[y' - (1/x)y = y^(-1)]*y y y' - (1/x)y^2 = 1 v = y^2, v'=2y y', y y' = v'/2 (1/2) v' - (1/x)v = 1 v' - (2/x)v = 2 e^(-2ln|x|) [v' - (2/x)v = 2] e^(-2ln|x|)v = 2 int [e^(-2ln|x|)] dx e^(-2ln|x|)v = 2 int [x^-2] dx e^(-2ln|x|)v = -2/x + C v = -2x^2/x + C x^2 v = -2x + C x^2; maybe but thats just me dabbling .....

OpenStudy (amistre64):

v = y^2 y^2 = -2x + C x^2 y = sqrt(-2x + Cx^2) y = sqrt(x(-2 + Cx)) y = sqrt(x) sqrt( Cx -2) which matches the wolfs